OpenStudy (dida):

A wire 24 inches long is to be cut into 4 pieces to form a rectangle whose shortest side has a length of x. a)Express the area A of the rectangle as a function of x.

OpenStudy (anonymous):

12x -x^2

OpenStudy (anonymous):

\[x(12-x)\] i think should work

OpenStudy (saifoo.khan):

the answer was a minute late, that should be in picoseconds!

OpenStudy (dida):

so both would work??

OpenStudy (anonymous):

idea is this if x is one side and y the other then the perimeter is \[2x+2y=24\] so \[x+y=12\] and \[y=12-x\] making the area (length times width) \[A(x)=x(12-x)\]

OpenStudy (anonymous):

yes they are the same \[A(x)=x(12-x)=12x-x^2\]

OpenStudy (dida):

ok and it says that i should make a conjecture about the dimensions that yeild a maximum area

OpenStudy (anonymous):

make a square (more or less obvious)

OpenStudy (anonymous):

i mean it is really obvious, because this is symmetric in x and y so it is biggest when they are equal. need some algebra to prove it?

OpenStudy (dida):

sure

OpenStudy (anonymous):

\[A(x)=12x-x^2\] a parabola that opens down. biggest at the vertex which is at \[-\frac{b}{2a}=-\frac{12}{-2\times-1}=6\] so x should be 6 and therefore y should also be 6 and so it is a square with area = 6^2=36

OpenStudy (anonymous):

hope you were not supposed to use calc or something

OpenStudy (dida):

yea this is the first few pages of calc

OpenStudy (dida):

what would be the domain?

OpenStudy (anonymous):

well you could do the following ridiculous thing 1) \[A(x)=12x-x^2\] 2) \[A'(x)=12-2x\] set \[12-2x=0\] get \[x=6\] but it is alway \[-\frac{b}{2a}\] for a quadratic

OpenStudy (dida):

the domain if we were to graph that. because it asks us to graph also

OpenStudy (anonymous):

well you are making a rectangle right? so negative lengths wouldn't make sense

OpenStudy (anonymous):

and clearly x cannot be bigger than 24 because that is all you got

OpenStudy (dida):

ok......

OpenStudy (anonymous):

actually scratch that

OpenStudy (anonymous):

x is only one side, so it cannot be bigger than 12

OpenStudy (dida):

alright

OpenStudy (anonymous):

i would say domain is \[0<x<12\] and here is a nice picture http://www.wolframalpha.com/input/?i=y%3D12x-x^2+for+0%3Cx%3C12

OpenStudy (anonymous):

there now you must be done with this homework problem because we have beaten it to death

OpenStudy (dida):

lol yes thanks you were very helpful XD